OpenStudy (anonymous):

Solve the equation. x/5 + 9 = 4 A. 65 B. –25 C. 5 D. 20

OpenStudy (anonymous):

First get rid of the 9 by doing the opposite function. The 9 is added to x, so remove it by using the opposite function, subtracting 9 from both sides. x/5 + 9 - 9 = 4 - 9 Now do the subtracting. x/5 = -5 Next, get rid of the 5 in the denominator of (x/5) by doing the opposite function. x is divided by 5, so remove it by using the opposite function, multiplying both sides by 5. (x/5) * 5 = (-5) * 5 Do the multiplying. (I put parentheses around the x/5 so you'll remember that this entire fraction is being multiplied by 5. I put parentheses around the -5 so you won't lose the negative sign or confuse it with a minus sign.) x = -25 So, x is equal to -25!

OpenStudy (haseeb96):

First put 4 on the left hand side then the sign of 4 will be change into -.